OpenStudy (tystar123):

The solution to m + (-38) = 15 is 53. True False

OpenStudy (steve816):

Well, what is 38 + 15?

OpenStudy (jtm-pa):

does 53 + -18 or 53-18 = 15?

OpenStudy (jtm-pa):

if true the answer is true if not the answer is false

OpenStudy (jtm-pa):

I meant to write 53-38

KjSaif (kjsaif):

Add 38 to both sides. m−38+38=15+38 m=53

OpenStudy (jtm-pa):

the answer would be true because 53-38= 15

OpenStudy (tystar123):

true or false?

KjSaif (kjsaif):

so the answer would be..........................

OpenStudy (jtm-pa):

:\ we both just proved it to you in different ways... the answer is true

KjSaif (kjsaif):

true! and true @jtm-PA